A monopolist faces the demand curve Q= 60-p/2. The cost function is C=Q
1
Expert's answer
2021-07-12T08:17:02-0400

Q = 60 - P/2 or P = 120 - 2Q.

Monopolist profit-maximizing quantity is produced, when MR = MC.


"MR = TR' = (P\u00d7Q)' = (120Q - 2Q^{2})' = 120 - 4Q,"

"MC = C' = (Q^{2})' = 2Q."

120 - 4Q = 2Q,

Q = 20 units.

P = 120 - 2(20) = 80.

"Ed = -b\u00d7P\/Q = -0.5\u00d780\/20 = -2,"

so the demand is elastic.